OpenStudy (anonymous):

Hello! If anyone could help me with this question, I would greatly appreciate it! Find the center and radius of the circle that passes through the points J(4, 4), K(-3, -1), and L(2, -8). Thank you very much!

Directrix (directrix):

@catsarecool1995 If you are allowed to use technology in your work, you can use the equation generator for a circle given three points of the circle. Go to http://www.mathportal.org/calculators/analytic-geometry/circle-through-three-points.php and plug in the three points of the circle.

Directrix (directrix):

The circle will be given in (x=h)^2 + (y-k)^2 = r^2 form. You'll have to pick out the coordinates of the center and the radius. Post what you get, and I'll tell you if that is what I got. Okay?

OpenStudy (anonymous):

Hi Directrix! This is the answer I got for this question, and I was coming here to just check it! I used the formula for a circle (x - h)^2 + (y - k)^2 Where (h , k) is the center r is the radius And what I got as an answer was: Center at (3 , -2), radius of sqrt(37) Would that be correct? =)

Directrix (directrix):

The circle generator produced the attached circle equation which has the same center and radius as you posted. So, your work agrees with mine. I hope we are correct. :)

OpenStudy (anonymous):

Thank you so much! And that equation generator is amazing! Very handy. Thanks again.

Directrix (directrix):

Glad to help.